OpenStudy (anonymous):

Minya is 30 years younger than her mom, and the sum of their ages is 58. How old is Minya?

Parth (parthkohli):

Let's say that mom is \(m\) years old, then Minya will be \(m - 30\). We know:\[ m + (m - 30) = 58 \]Solve for \(m - 30\).

OpenStudy (anonymous):

I thought that to, but online it says minya is 14 but it doesnt explain how.

Parth (parthkohli):

I'd explain that.

Parth (parthkohli):

can you first solve the equation?

OpenStudy (anonymous):

Okay if it is m-30 = 58 then you would do the inverse of it which would be 58 - 30 = 28.

Parth (parthkohli):

?

OpenStudy (anonymous):

this is an equation.

Parth (parthkohli):

\[ m - 30 = 58\]\[ m = 58 + 30\]

OpenStudy (anonymous):

58 - 30 = 28

OpenStudy (anonymous):

oh sorry 58 + 30 =88

Parth (parthkohli):

Yeah

Parth (parthkohli):

\[2m = 88 \implies m = 44 \implies m - 30 = 14 \]